​(I) Four point charges are located at the vertices of a square with side L.


We give q2 ​= q3 ​= Q and q1​= q4 ​= −2Q.


​Determine the resulting electric field​ (a) at point A​,


in the center of the square; (b) at point B


1
Expert's answer
2022-02-13T17:45:19-0500

"\vec E=\vec E_1+\vec E_2+\vec E_3+\vec E_4,"

a)

"E=2\sqrt 2\frac {kq}{L^2},"

b)

"E_x=\frac{8kq}{5\sqrt 5 L^2},"

"E_y=\frac{4kq(5\sqrt 5 -3)}{5\sqrt 5 L^2},"

"E=\sqrt{E_x^2+E_y^2}=\frac{4kq(138-30\sqrt 5)}{5\sqrt 5L^2}."